x^2+5x=?? find the number that should be added to each expression to make it a perfect square thanks for the help!
Answer by lwsshak3(11628)

x^2+5x=?? find the number that should be added to each expression to make it a perfect square
Take half of the coefficient of the x-term=5/2
Square this number=25/4
This is the number to add to make the expression a perfect square=x^2+5x+25/4
To keep the equation in balance you must also subtract this number from the perfect square
like,(x^2+5x+25/4)-25/4=(x+5/2)^2-25/4